### v1

@@ 3.5.1 Algorithm Selection
To establish a relationship among the bug types, information categories, and the key editing actions necessary for their reproduction, we use the Apriori apriori algorithm. Apriori is a well-known algorithm for mining frequent item sets from a list of transactions. It helps one identify common patterns and associations between different elements. In our context, we apply the Apriori algorithm to analyze the information gathered during bug reproduction. Our goal was to determine the frequent combinations of bug types, information categories, and edit actions during bug reproduction. We leveraged the Apriori algorithm’s systematic pattern-mining capabilities to establish these relationships.
@@ 3.5.4 Association Rule Generation
We conducted two separate association rule mining operations in our study - one focused on information categories while the other focused on edit actions used to reproduce deep learning bugs. In particular, we extracted 27 itemsets and 34 rules for information categories, highlighting the useful information for reproducing deep learning bugs. Similarly, we extracted 126 item sets and 284 rules for edit actions, capturing the edit actions to reproduce bugs. We generated association rules based on the entries in our dataset and did not filter or remove any rules before determining confidence values. We then calculate the confidence values for all generated rules to identify the most influential ones for connecting bug types with edit actions and useful information.
@@ 3.5.5 Computation of Confidence Values for the Generated Association Rules
As discussed earlier, confidence indicates how frequently a rule occurs, while support indicates the generality of the rule. To compute the confidence values for each association rule, we performed the following steps:
@@ 3.6 User Study
To assess the benefits and implications of our findings in a real-life setting, we conduct a user study involving 22 users (10 from academia + 12 from industry) (Step 6, Fig. 1). We discuss our study setup, including instrument design and participation selection, as follows:
@@
Introduction: We first summarize our findings on the reproducibility of deep learning bugs to provide the developers with the necessary context and background information.
@@
Questionnaire Preparation: First, we select eight bugs (2 Tensor, 2 API, 2 Model, and 2 Training Bugs) from our dataset. Second, we prepare four sets of 2 bugs each by randomly selecting bugs from those eight bugs. Third, we provide the users with the issue description and the code snippet from the original Stack Overflow post. We also provide a Google Colaboratory notebook containing the code for sample edit operations to aid the bug reproduction process. Finally, we include a free-text box in the form to allow the users to share any additional information about edit actions not covered in our study.
Study Session: During our study session, each participant completes the following five tasks. First, the participant provides their demographic information. Second, the participant explains their daily challenges when reproducing deep learning bugs. Third, the participant reproduces two deep learning bugs and reports the edit actions and information they used to reproduce the bugs. Fourth, the participant also provides the rationale behind their used edit actions and information categories to reproduce the bugs. Finally, the participant provides information about any other edit actions that they might have used to reproduce the bugs but are not covered in our study.
@@
Participant Selection: We first conducted a pilot study with two researchers and two developers. Based on their feedback, we rephrased ambiguous questions and added sample code to aid the manual bug reproduction by the users. Incorporating this constructive feedback enabled us to refine and improve the quality of our final questionnaire. Then, we invite professional developers and researchers with relevant deep learning experience to our study. We send our invitations to the potential participants using direct correspondences, organization mailing lists (e.g., Mozilla Firefox), and public forums (e.g., LinkedIn and Twitter). A total of 22 participants responded to our invitations. Out of them, 10 (45%) came from academia, and 12 (55%) came from industry. In terms of bug-fixing experience with deep learning bugs, 14 participants (63.63%) had 1-5 years of experience, 4 (18.18%) had 5-10 years of experience, and the remaining 4 (18.18%) had less than one year of experience. In terms of deep learning frameworks, 19 participants (86.34%) reported having working experience with Tensorflow, 21 (95.45%) reported experience with PyTorch, and 20 (90.91%) reported experience with Keras. All these statistics indicate a high level of cross-framework expertise within our participants. We carefully divided the developers into control and experiment groups, as shown in Table. 4. Our guiding principle was to ensure that both groups had a similar distribution in their relevant experience. Specifically, we surveyed all the developers about their experience and used a stratified random sampling approach to assign developers to the two groups. This randomization allowed us to minimize potential bias and confounding factors across the groups. Using these two groups, we wanted to assess the benefit of our recommended information in the context of bug reproduction. The control group receives no hints about how to reproduce the bug. In contrast, the experimental group receives hints (e.g., useful information, edit actions) that could help them reproduce a bug.
@@
| Training (T) | 26 | CNN, LSTM, AutoEncoder, MLP, RCNN, |
@@
| Model (M) | 20 | BERT, CNN, GMM, LSTM, MLP |
@@
| Tensor and Input (I) | 21 | CNN, GAN, Logistic Regression, MLP, ResNet |
@@
3.7 User Study Results Analysis
We perform an agreement analysis between our empirical findings and the findings from the user study to assess their generalizability and strengthen our understanding of the reproducibility of deep learning bugs. To perform the agreement analysis, we compare the edit actions used in our manual bug reproduction process with those used by participants in the user study. If similar edit actions were used, it would indicate a strong agreement between the empirical findings and user study results.
In our user study, the participants first reproduce their assigned bugs and report their used information or actions to reproduce their bugs. To analyze the effectiveness of our recommended information, we compare the control and experimental groups in terms of their success rates in bug reproduction and the time taken to reproduce the bugs. We select bug reproducibility rate and time taken for bug reproduction as our key metrics to evaluate the effectiveness of our findings. These quantitative metrics directly measure how successful and efficient our recommended information is in assisting the developers to reproduce deep learning bugs. Higher reproducibility rates and reduced reproduction times in the experimental group compared to the control group would indicate that our findings are effective for improving deep learning bug reproducibility. We also analyze the qualitative feedback from the participants to determine if provided hints were useful or not. We also analyze them to uncover new insights into deep learning bug and their reproduction.
@@ 4.1 RQ1: Which edit actions are crucial for reproducing deep learning bugs?
We reproduced 85 bugs of different types and architectures to answer the first research question. Table. 5 briefly summarizes our reproduced bugs. Through our comprehensive reproduction process, we identify ten edit actions that are crucial for reproducing deep learning bugs. Table. 3 shows the identified actions from our qualitative analysis. We explain these actions as follows.
Input Data Generation (A) is one of the key edit actions for reproducing deep learning bugs. This action involves generating input data that simulates the data used for training. We generate the input data using the data shape and distribution provided by the reporter in a bug report. Our manual bug reproduction shows that 73% of our collected posts from Stack Overflow have the relevant data characteristics. For example, the following text from the Stack Overflow post (Issue #61781193) shows how the reporter might submit the input data distribution.
@@
Neural Network Construction (A) was one of the most used edit actions during our bug reproduction. In this edit action, we construct a neural network based on the architecture provided by the reporter. Similar to the data characteristics, the information about the neural network is present in 65% of the reproducible issue reports. Using the neural network description from the issue reports, we were able to construct the models. For example, the following text from the Stack Overflow post (Issue #63204176) mentions the neural network but does not provide a code snippet for the network.
@@
Using this partial information, we constructed multiple multinomial logistic regression models, and despite the lack of relevant code examples, we successfully reproduced the corresponding bug.
@@
Hyperparameter Initialization (A) is one of the core edit actions for reproducing deep learning bugs. As a part of this edit action, we initialize various hyperparameters (e.g., number of epochs, batch size, optimizer) for training the neural networks. Sometimes, the reporter did not provide these configurations, and we had to initialize them using default parameters to reproduce the bug. This is further evidenced by the fact that only 53% of our collected posts from Stack Overflow have information about the hyperparameters. For example, the following code snippet from the Stack Overflow post (Issue #31880720) shows a case where the hyperparameters were not initialized.
@@
To reproduce the bug, we initialized the batch size with commonly-used values and the number of epochs ranging from 1 to 10. Since we initialized these hyperparameters, we ran the edited code snippet five times to confirm the effectiveness of our edit operation. We were able to reproduce the bug in all five iterations. We performed the steps mentioned above for all 37 bugs involving hyperparameter initialization.
@@
Import Addition and Dependency Resolution (A) are one of the most common edit actions for all types of bugs. In this edit action, we analyze the code manually and determine the dependencies required for the code snippet to run. Then, we install the dependencies and manually import them to complete the code snippet. This edit action was also used to reproduce traditional software bugs, as reported by Mondal et al. msr2019mondal . In the Stack Overflow posts we collected, 47% of the issue reports lacked the required dependency information. This significant absence of the dependency information further highlights the need for our suggested edit action. For instance, the code snippet reported in the Stack Overflow post (Issue #50306988) was incomplete, and the dependency details were missing. Hence, we resolved the dependencies (keras, numpy and random) and added the required import statements to the code snippet. With the resolved dependencies, we were able to reproduce the bug successfully.
@@
Obsolete Parameter Removal (A) is a vital edit action to enhance the reproducibility of bugs where the differences in the library and framework versions may lead to problems. In this edit action, we remove all the obsolete parameters not supported by the newer version of the library or framework. For example, the code snippet in the Stack Overflow post (Issue #65992364) has the following line of code.
@@
Compiler Error Resolution (A) is one of the extensively employed edit actions in reproducing bugs. In this edit action, we resolve the compiler errors to reproduce the bug. For example, in the Stack Overflow post (Issue #71514447), the reporter reports that the training loss is significantly increasing after every epoch. They suspect that the bug might be due to the incorrect computation of loss values. Furthermore, to help the developers reproduce the bug, they provide a detailed description of the bug and a complementary code snippet. However, the code snippet provided by the user does not compile, as the criterion function has not been defined. Hence, we replaced the function parameter with the default value of cross-entropy loss, as shown below.
@@
Dataset Procurement (A) is a critical edit action to reproduce bugs that require specific datasets. In this edit action, we analyze the issue report and attempt to procure the dataset mentioned in the report. For instance, in the Stack Overflow post (Issue #73966797), the reporter mentioned that they used the CIFAR-10, a well-known dataset for object detection. To reproduce the bug, we downloaded the dataset from its original source666https://www.cs.toronto.edu/ kriz/cifar.html, and using the dataset and the training code, we could reproduce the bug.
Downloading Models & Tokenizers (A) is one of the edit actions that was frequently used to reproduce bugs in Large Language Models and Transformer-based architectures. In this edit action, we download the pre-trained models and tokenizers for reproducing a bug. For example, in the Stack Overflow post (Issue #69660201), the reporter has provided the following information.
@@
Version Migration (A) is one of the commonly used edit actions for all types of bugs. In this edit action, we replace the outdated or deprecated versions of a library or framework with the latest version to reproduce the bug. Mondal et al. msr2019mondal introduced a similar edit operation known as ‘Code Migration’ in their work. This operation often confirms the validity of the bug by reproducing it with the latest version. For example, the post on Stack Overflow (Issue #45711636) describes an issue with the CNN architecture constructed by the user. However, the code snippet provided in the issue uses Tensorflow 1.3.0. To reproduce this issue in TensorFlow 2.14.0, we carefully migrated the model built in TensorFlow 1.3.0 to the syntax of Tensorflow version 2.14.0. Specifically, we updated the Sequential, Conv2D and MaxPooling2D constructors to match the TensorFlow 2.14.0 API syntax. We also changed padding and pool size parameter schemes and removed the obsolete input shapes. Finally, we upgraded the model compilation and imported the required Tensorflow 2.14.0 modules. After modifying the code snippet, we could reproduce the bug successfully in our runtime environment.
4.2 RQ2: What information categories and edit actions are useful for reproducing specific types of deep learning bugs?

### v2

@@ 4.6.1 Algorithm Selection
To establish a relationship among the bug types, component information, and the key editing actions necessary for bug reproduction, we use the Apriori apriori algorithm. Apriori is a well-known algorithm for mining frequent itemsets from a list of transactions. It helps one identify common patterns and associations between different elements.
The Apriori algorithm exploits the principle that if an itemset is frequent across the transactions, then all of its subsets must also be frequent. It starts by identifying frequent individual items in the dataset and extends them to larger and larger itemsets as long as they meet certain constraints (e.g., support threshold). The algorithm terminates when no further successful extensions are found. More specifically, the Apriori algorithm consists of two main steps: the join step and the prune step. In the join step, the algorithm generates new candidate itemsets by joining the frequent itemsets found in the previous iteration. In the prune step, the algorithm checks the support count of each candidate itemset and discards the itemsets that do not meet the minimum support threshold. This process is repeated until no more frequent itemsets are generated. In our context, we apply the Apriori algorithm to analyze the information gathered during bug reproduction. Our goal was to determine the frequent combinations of bug types and component information, as well as edit actions during bug reproduction. We leveraged the Apriori algorithm’s systematic pattern-mining capabilities to establish these relationships. By identifying the frequent itemsets, we can gain insights into the common patterns and associations among bug types, component information, and edit actions, which can help us understand and improve the bug reproduction process.
@@ 4.6.3 Metrics for Apriori Algorithm
4.6.4 Definitions for Apriori Algorithm
Itemset: An itemset is a set of one or more items. In our context, an item can be a bug type, an information category, or an edit action. For example, {T, D, H} is an itemset containing three items: bug type T (training bug), information category D (dataset), and information category H (hyperparameters).
Transaction: A transaction is a record that contains one or more itemsets. In our study, we have two types of transactions as follows:
Bug Type Information Category: These transactions associate a bug type with the useful component information for reproducing that bug.
Bug Type Edit Action: These transactions associate a bug type with the edit actions performed to reproduce that bug.
Rule: A rule is an implication of the form , where and are itemsets. It suggests that if itemset is present in a transaction, then itemset is likely to be present as well. In our study, rules are generated from the transactions to establish associations between bug types and component information or edit actions.
@@ 4.6.5 Association Rule Generation
We conducted two separate association rule mining operations in our study - one focused on component information while the other focused on edit actions used to reproduce deep learning bugs. The Apriori algorithm generates rules by first identifying frequent itemsets and then creating rules from them. The steps below explain the process of rule generation with an example.
Identify frequent itemsets: The algorithm scans the transactions to find itemsets that occur frequently while satisfying the minimum support threshold. For example, if the item T, D appears in 20% of the transactions and the minimum support threshold is 10%, it is considered a frequent item.
Generate rules: Once frequent itemsets are identified, the algorithm generates rules from them. For each frequent itemset, the algorithm creates rules by splitting the itemset into antecedent (left-hand side) and consequent (right-hand side). For example, from the item T, D, H, the following rules can be generated:
Calculate confidence: For each generated rule, the algorithm calculates the confidence value. Confidence measures the likelihood that the consequent itemset appears in a transaction given that the antecedent itemset is present. Rules with confidence values above a minimum threshold are considered strong associations.
In particular, we extracted 27 itemsets and 34 rules for component information, highlighting the useful information for reproducing deep learning bugs. Similarly, we extracted 126 itemsets and 284 rules for edit actions, capturing the edit actions needed to reproduce bugs. We generated association rules based on the entries in our dataset and did not filter or remove any rules before determining confidence values. We then calculate the confidence values for all generated rules to identify the most influential ones for connecting bug types with edit actions and useful information.
@@ 4.6.6 Computation of Confidence Values for the Generated Association Rules
As discussed earlier, support indicates how frequently a rule occurs, while confidence indicates the generality of the rule. To compute the confidence values for each association rule, we performed the following steps:
@@ 4.7 User Study
To assess the benefits and implications of our findings in a real-life setting, we conduct a user study involving 22 developers (10 from academia + 12 from industry) (Step 6, Fig. 2). We discuss our study setup, including instrument design and participation selection, as follows:
@@
Introduction: We first summarize our findings on the reproducibility of deep learning bugs to provide the participants with the necessary context and background information. For the survey itself, we do not give the respondents a fixed time to complete it, but we specify that it should take 60 minutes on average; this number was derived from our pilot study. This was done to ensure that the respondents do not work under time pressure. Since the participants have different levels of experience, allocating a fixed time for bug reproduction might affect our results.
@@
Questionnaire Preparation: First, we select eight bugs (2 Tensor, 2 API, 2 Model, and 2 Training Bugs) from our dataset constructed during manual bug reproduction and dataset creation. During this process, we categorized the bugs by difficulty level and type based on the number of edit actions and critical information required to reproduce them. We use stratified random sampling to pick 2 bugs from each type; one of the bug types is relatively easy to reproduce with only 1 edit action, and the other type is relatively difficult to reproduce warranting multiple edit actions. We pick 8 bugs following this approach and, then randomly assign them to four sets, each containing 1 easy and 1 difficult bug. Second, we provide the users with the issue description and the code snippet from the original Stack Overflow post. We also provide a Google Colaboratory notebook containing the code for sample edit operations to aid the bug reproduction process. Finally, we include a free-text box in the form to allow users to share any additional information about edit actions not covered in our study.
Study Session: During our study session, each participant completes the following five tasks. First, the participant provides their demographic information. Second, the participant explains their daily challenges when reproducing deep learning bugs. Third, the participant reproduces two deep learning bugs and reports the edit actions and information they used to reproduce the bugs. Fourth, the participant also provides the rationale behind their used edit actions and component information to reproduce the bugs. Finally, the participant provides information about any other edit actions that they might have used to reproduce the bugs but are not covered in our study.
@@
Participant Selection: We first conducted a pilot study with two researchers and two developers. Based on their feedback, we rephrased ambiguous questions and added sample code to aid the manual bug reproduction by the users. Incorporating this constructive feedback enabled us to refine and improve the quality of our final questionnaire. Then, we invite professional developers and researchers with relevant deep learning experience to our study. We send our invitations to the potential participants using direct correspondences, organization mailing lists (e.g., Mozilla Firefox), and public forums (e.g., LinkedIn and Twitter). A total of 22 participants responded to our invitations. Out of them, 10 (45%) came from academia, and 12 (55%) came from industry. In terms of bug-fixing experience with deep learning bugs, 14 participants (63.63%) had 1-5 years of experience, 4 (18.18%) had 5-10 years of experience, and the remaining 4 (18.18%) had less than one year of experience. In terms of deep learning frameworks, 19 participants (86.34%) reported having working experience with Tensorflow, 21 (95.45%) reported experience with PyTorch, and 20 (90.91%) reported experience with Keras. All these statistics indicate a high level of cross-framework expertise within our participants.
Defining Control and Experimental Groups: We carefully divided the study participants into control and experiment groups, as shown in Table. 4. Using these two groups, we wanted to assess the benefit of our recommended information in the context of bug reproduction. The control group receives no hints about how to reproduce a bug. In contrast, the experimental group receives hints (e.g., useful information, edit actions) that could help them reproduce a bug.
Ensuring Similar Experience Levels: Our guiding principle was to ensure that both groups had a similar distribution in their relevant experience. Specifically, we surveyed all the developers about their experience and used a stratified random sampling approach to assign them to the two groups. This randomization allowed us to minimize potential bias and confounding factors across the groups.
Leveraging Associations to Produce Hints for User Study: In Step 3.6.6, we derived high-confidence associations to develop hints for the user study. These associations revealed edit actions and critical information needed to reproduce specific bugs. We then use this information to prepare the hints for the user study. In the user study, we do not directly provide the solution for reproducing the bugs. Instead, we give them a small hint (i.e., where to look in the bug report for one of the pieces of critical information or how to start the reproduction process), which helps them in bug reproduction. With this intervention, we plan to measure if our findings can help a person reproduce the bug more accurately or quickly.
4.8 User Study Results Analysis
To assess the effectiveness of our edit actions, we analyze the participants’ responses and qualitative feedback. Specifically, we analyze the edit actions used by the participants and compare them with those used in our manual bug reproduction process. If similar edit actions were used, it would indicate that our recommended edit actions and component information were effective in the reproduction of the bugs.
In our user study, the participants first reproduce their assigned bugs and report their used information or actions to reproduce their bugs. To analyze the effectiveness of our recommended information, we compare the control and experimental groups in terms of their success rates in bug reproduction and the time taken to reproduce the bugs. We select bug reproducibility rate and time taken for bug reproduction as our key metrics to evaluate the effectiveness of our findings. These quantitative metrics directly measure how successful and efficient our recommended information is in assisting the developers to reproduce deep learning bugs. Higher reproducibility rates and reduced reproduction times in the experimental group compared to the control group would indicate that our findings are effective for improving deep learning bug reproducibility. We also analyze the qualitative feedback from the participants to determine if our provided hints were useful or not. We also analyze them to uncover new insights into deep learning bug and their reproduction.
@@ 4.8 User Study Results Analysis
| Training (T) | 50 | CNN, LSTM, AutoEncoder, MLP, RCNN, |
@@
| Model (M) | 42 | BERT, CNN, GMM, LSTM, MLP |
@@
| Tensor and Input (I) | 29 | CNN, GAN, Logistic Regression, MLP, ResNet |
@@ 5.1 RQ1: Which edit actions are crucial for reproducing deep learning bugs?
To answer the first research question, we worked with 165 bugs, and reproduced 148 bugs of different types and architectures. Table. 5 briefly summarizes our reproduced bugs. Through our comprehensive reproduction process, we identify ten edit actions that are crucial for reproducing deep learning bugs. Table. 6 shows the identified actions from our qualitative analysis. We explain these actions as follows.
Input Data Generation (A1) is one of the key edit actions for reproducing deep learning bugs. This action involves programmatically generating synthetic input data that closely matches the characteristics of the original data used for training the model. The key objective of input data generation is to simulate representative data that can trigger or reproduce the erroneous model behavior described in the bug report. This allows for the reproduction of issues that manifest only in the presence of specific data properties or distributions. To perform input data generation, we leverage any details about the data that are provided in the bug report, such as data types, value ranges, shapes, distributions, preprocessing steps, etc. For example, for image data, the report may specify that inputs are RGB images of size 224x224x3 with pixel values normalized to [0,1]. Similarly, for text data, the description may indicate sequences of 512 tokens processed using a particular tokenizer. Using this information, we can systematically generate synthetic data matching the properties through appropriate library functions. For images, we can use libraries like OpenCV OpenCV2024 or PIL PIL to construct random images of the required size and channels. For text, we can sample token sequences from a standard corpus or use specialized generative models like GPT-2 radford2019language . Our manual bug reproduction shows that 73% of our collected posts from Stack Overflow have the relevant data characteristics. Let us consider the issue reported in the Stack Overflow post (Issue #61781193). In this post, the reporter suspects that the model is not learning - as evidenced by the constant training loss across the epochs. The following text from the post shows how the reporter might submit the input data distribution.
@@
Neural Network Construction (A2) was one of the most used edit actions during our bug reproduction. In this edit action, we construct a neural network based on the architecture provided by the reporter. Similar to the data characteristics, the information about the neural network is present in 65% of the reproducible issue reports. Using the neural network description from the issue reports, we were able to construct the models. Let us consider the issue reported in the Stack Overflow post (Issue #63204176). In this post, the reporter submits an issue where a CrossEntropy loss function within a loop is overwritten with a Tensor, causing a TypeError in later iterations. When we analyze the post, we find that the reporter mentions that they have used a logistic regression model (1-layer neural network with a sigmoid activation function logisticRegression ). However, the reporter does not provide the code snippet necessary for reproducing the bug.
@@
To reproduce the bug, we constructed multiple logistic regression models for the MNIST dataset, as highlighted in the post. We added the import statements, wrote the code to load the MNIST dataset, and completed the code. Using the training loop’s code snippet from the original code and our edit actions, we could complete the code snippet and reproduce the bug successfully. Using this partial information, we constructed multiple multinomial logistic regression models, and despite the lack of relevant code examples, we successfully reproduced the corresponding bug.
| Edit Action | Overview |
| A1: Input Data Generation (D) | Generating input data that simulates the data used for training the model. |
| A2: Neural Network Construction (N) | Reconstructing or modifying the neural network based on the information provided. |
| A3: Hyperparameter Initialization (H) | Initializing the hyperparameters for training, such as batch size and number of epochs. |
| A4: Import Addition and Dependency Resolution (R) | Determining the dependencies in the code snippet and adding the missing import statements. |
| A5: Logging (L) | Adding appropriate logging statements to capture relevant information during reproduction. |
| A6: Obsolete Parameter Removal (O) | Removing outdated parameters or functions to match the parameters of the latest library versions. |
| A7: Compiler Error Resolution (C) | Debugging and resolving compiler errors that arise due to syntactic errors in the provided code snippet. |
| A8: Dataset Procurement (P) | Acquiring the necessary datasets and using them to train the model. |
| A9: Downloading Models & Tokenizers (M) | Fetching pre-trained models and tokenizers from external sources. |
| A10: Version Migration (V) | Updating code to adapt to changes introduced in newer versions of libraries and frameworks. |
@@
Hyperparameter Initialization (A3) is one of the core edit actions for reproducing deep learning bugs. As a part of this edit action, we initialize various hyperparameters (e.g., number of epochs, batch size, optimizer) for training the neural networks. Sometimes, the reporter did not provide these configurations, and we had to initialize them using default parameters to reproduce the bug. This is further evidenced by the fact that only 53% of our collected posts from Stack Overflow have information about the hyperparameters. Let us consider the issue reported in the Stack Overflow post (Issue #31880720), where the author gets a poor test accuracy of 13.9% when training a neural network on a synthetic binary classification dataset. After manual analysis of the post and the code snippet, we observe that some of the hyperparameters in the code snippet have not been initialized.
@@
To reproduce the bug, we initialized the batch size with commonly-used values and the number of epochs ranging from 1 to 10. Since we initialized these hyperparameters, we ran the edited code snippet five times to confirm the effectiveness of our edit operation. We were able to reproduce the bug in all five iterations. We performed the steps mentioned above for all 56 bugs involving hyperparameter initialization.
@@
Import Addition and Dependency Resolution (A4) are one of the most common edit actions for all types of bugs. In this edit action, we analyze the code manually and determine the dependencies required for the code snippet to run. Then, we install the dependencies and manually import them to complete the code snippet. This edit action was also used to reproduce traditional software bugs, as reported by Mondal et al. msr2019mondal . In the Stack Overflow posts we collected, 47% of the issue reports lacked the required dependency information. This significant absence of the dependency information further highlights the need for our suggested edit action. Let us consider the issue reported in the Stack Overflow post (Issue #50306988). In this post, the neural network model with softmax activation struggles to fit a simple 2-feature classification dataset, converging extremely slowly compared to logistic regression, which achieves 100% accuracy. However, the code snippet reported in the post was incomplete, and the dependency details were missing. Hence, we resolved the dependencies (keras, numpy and random) and added the required import statements to the code snippet. With the resolved dependencies, we were able to reproduce the bug successfully.
@@
Obsolete Parameter Removal (A6) is a vital edit action to enhance the reproducibility of bugs where the differences in the library and framework versions may lead to problems. In this edit action, we remove all the obsolete parameters not supported by the newer version of the library or framework. Consider the issue reported in the Stack Overflow post (Issue #65992364). In this post, the reporter attempts to optimize an object detection model using ‘pytorch-mobile’, but the code snippet fails to optimize the model file size. The code snippet in the issue also contains the following line of code.
@@
Compiler Error Resolution (A7) is one of the extensively employed edit actions in reproducing bugs. In this edit action, we resolve compiler errors to get the code running and reproduce the bug. While downgrading the compiler or library versions can sometimes resolve errors or enable the bug reproduction with deprecated functionality, it may not always be feasible. For example, major framework releases often remove support for older versions, making downgrading impossible. Additionally, Python version mismatches between the original buggy code and the current development environment can prevent successful downgrading. In cases where downgrading is infeasible due to such constraints, correction of compiler errors enables reproducing deep learning bugs despite incompatible environments.
For example, in the Stack Overflow post (Issue #71514447), the reporter states that the training loss is significantly increasing after every epoch. They suspect that the bug might be due to the incorrect computation of loss values. Furthermore, to help the developers reproduce the bug, they provide a detailed description of the bug and a complementary code snippet. However, the code snippet provided by the user does not compile, as the criterion function has not been defined. Hence, we replaced the function parameter with the default value of cross-entropy loss, as shown below.
@@
Dataset Procurement (A8) is a critical edit action to reproduce bugs that require specific datasets. In this edit action, we analyze the issue report and attempt to procure the dataset mentioned in the report. For instance, in the Stack Overflow post (Issue #73966797), the reporter mentioned that they used the CIFAR-10, a well-known dataset for object detection. To reproduce the bug, we downloaded the dataset from its original source666https://www.cs.toronto.edu/ kriz/cifar.html, and using the dataset and the training code, we could reproduce the bug.
Downloading Models & Tokenizers (A9) is one of the edit actions that was frequently used to reproduce bugs in Large Language Models and Transformer-based architectures. In this edit action, we download the pre-trained models and tokenizers for reproducing a bug. Let us consider the issue reported in the Stack Overflow post (Issue #69660201). In this post, the reporter faces a ValueError when fitting a text classification model with a BERT tokenizer, due to a mismatch between the model’s expected input and the tf.data.Dataset created from the text corpus and labels. For example, in the Stack Overflow post (Issue #69660201), the reporter has provided the following information.
@@
Version Migration (A10) is one of the commonly used edit actions for all types of bugs. In this edit action, we replace the outdated or deprecated versions of a library or framework with the latest version to reproduce the bug. Mondal et al. msr2019mondal introduced a similar edit operation known as ‘Code Migration’ in their work. This operation often confirms the validity of the bug by reproducing it with the latest version. While downgrading the library to an older version can help reproduce bugs relying on deprecated functionality, upgrading to a more recent version is still useful when downgrading is impossible due to dependency conflicts. Furthermore, upgrading to the recent version can also help determine if the bug persists across different versions, providing valuable information for developers. Moreover, when the developers set up the environment for the bug reproduction, installing the latest compatible versions of libraries and frameworks can be more straightforward than attempting to match the exact outdated versions mentioned in the bug report. In cases where downgrading is infeasible, upgrading to the recent version can help reproduce deep learning bugs. For example, the post on Stack Overflow (Issue #45711636) describes an issue with the CNN architecture constructed by the user. When the user passes the input through the CNN, they encounter a ValueError, which highlights a problem with the negative dimension value of the input. However, the code snippet provided in the issue uses Tensorflow 1.3.0. To reproduce this issue in TensorFlow 2.14.0, we carefully migrated the model built in TensorFlow 1.3.0 to the syntax of Tensorflow version 2.14.0. Specifically, we updated the Sequential, Conv2D and MaxPooling2D constructors to match the TensorFlow 2.14.0 API syntax. We also changed padding and pool size parameter schemes and removed the obsolete input shapes. Finally, we upgraded the model compilation and imported the required Tensorflow 2.14.0 modules. After modifying the code snippet, we could reproduce the bug successfully in our runtime environment.
Summary of RQ1: By manually reproducing 148 deep learning bugs, we identify ten key edit actions that could be useful to reproduce deep learning bugs (e.g., input data generation, neural network construction, hyperparameter initialization). These edit actions can help developers complete the code snippets and thus reproduce their deep learning bugs.
5.2 RQ2: What component information and edit actions are useful for reproducing specific types of deep learning bugs?